OpenStudy (anonymous):

what is the solution of 3(2x-1)- 2(3x+4)=11x

OpenStudy (king):

3(2x-1)=6x-3 -2(3x+4)=-6x-8 so adding, -11 11x=-11 x=-1

OpenStudy (anonymous):

thank you